OpenStudy (anonymous):

Find the midpoint of the line segment. Point 1- (5,6) Point 2 - (-3,-2)

hero (hero):

(1,2)

OpenStudy (anonymous):

x=5-(-3)=8; y=6-(-2)=8. midpoint (-3+4,-2+4)=(1,2)

myininaya (myininaya):

this is it you guys first girl to become a hundred person! :) \[M=(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})=(\frac{5+(-3)}{2},\frac{6+(-2)}{2})=(\frac{2}{2},\frac{4}{2})\]

myininaya (myininaya):

giving you what hero has

myininaya (myininaya):

and wcolmen

hero (hero):

I did it graphically

hero (hero):

I hate using that formula

hero (hero):

If a problem can be solved geometrically rather than algebraically, I prefer to use the geometric method

OpenStudy (anonymous):

Thanks Guys!!